3 A of current is flowing in a long straight wire. A positive charge, q = 7 C is
moving parallel to the wire at a distance of 0.05 m with a velocity of 250 ms-1.
Calculate the magnitude of the magnetic field due to the flow of current in the
wire.
1
Expert's answer
2020-06-03T12:12:47-0400

As per the given question,

current (I)=3A

positive charge (q)=7c

distance(r) =0.05 m

velocity (v) = 250 m/sec

magnitude of the magnetic field (B)=?

Magnitude of the magnetic field due to the flow of the current in the wire

"(B)=\frac{\mu_o I}{2 \pi r}"

Now, substituting the values in the question,

"(B)=\frac{10^{-7} \times 3}{0.05}=0.6\times 10^{-5}T"